OpenStudy (anonymous):

The ratio of the numerator to the denominator of a fraction is 2 to 3. If both the numerator and the denominator are increased by 2, the fraction becomes 3/4. What is the original fraction? Which of the following systems of equations can be used to solve the problem? n + 4 = 3 and d + 5 = 4 3n - 2d = 0 and 4n + 2 = 3d + 2 3n = 2d and 4n + 8 = 3d + 6

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow {2x + 2 \over 3x + 3} = {3 \over 4} }\) \(\Large \color{purple}{\rightarrow 4(2x + 2) = 3(3x + 3) }\) \(\Large \color{purple}{\rightarrow 8x + 8 = 9x + 9}\) \(\Large \color{purple}{\rightarrow 8x = 9x + 1 }\) \(\Large \color{purple}{\rightarrow x = -1 }\)

OpenStudy (anonymous):

So you need 99 in engagement in order to graduate from high school?

OpenStudy (anonymous):

ParthKohli has the correct method. Answer to the question is 3n - 2d = 0 and 4n + 2 = 3d + 2